Two small plastic spheres are given positive electrical charges. When they are a distance of 15.1 cm apart, the repulsive force between them has a magnitude of 0.230 N .

1) What is the charge on the first sphere if it equals one-quarter the charge of the second sphere?

2) What is the charge on the second sphere described in the previous part?
Answer #1

Given that,

q1 = 0.25 q2

or

q2 = 4 q1

Now as per coulomb's law of electrostatics

F = k q1 q2 / r^2

0.23 = 9*10^9* 4* q1^2 / 0.151^2

q1 = 3.817*10^-7 C

=============

b)

q2 = 4 q1 = 1.527*10^-6 C

============

do comment in case any doubt, will reply for sure Goodluck